OpenStudy (anonymous):

solve 7-(-y-5)=2(y+3)-6(y+3).. need to know steps on how to get the answer

OpenStudy (gw2011):

7-(-y-5)=2(y+3)-6(y+3) 7+y+5=2y+6-6y-18 y-2y+6y=6-18-7-5 5y=-24 y=-24/5=-4 4/5

OpenStudy (anonymous):

looks so confusing =/ how did you get y-2y+6y=6-18-7-5

jhonyy9 (jhonyy9):

those terms with y take on left part of = and those terms without y on oder part of equal sign

jhonyy9 (jhonyy9):

there you have 7+y+5=2y+6-6y-18 so those terms with y take on the left part of = and will get but before all you need to know when will be changed the place of one term from right on left part or invers this term will changed the sign for example from (-y) will be +y or invers from +y will be -y

jhonyy9 (jhonyy9):

so if we know this property from this result without parantheses we will get y-2y+6y=6-18-7-5 -hope that this is understandably

jhonyy9 (jhonyy9):

- this ALL you will can get it

jhonyy9 (jhonyy9):

so this ALL you will can get it if you will add to both sides

jhonyy9 (jhonyy9):

terms with different signs (-) or (+) depended on how you can elimine those terms with or without y for you arrange those all in one part of equal sign those terms with y and the oder side without y

jhonyy9 (jhonyy9):

depended how you have learned in your schoole,how is it in your mathbook

OpenStudy (anonymous):

thanks =)

jhonyy9 (jhonyy9):

ok my pleasure